OpenStudy (anonymous):

a=bh a=72 h=9 substitue the given values into the formula an solve for the unknown valuble

OpenStudy (anonymous):

72 = b * 9 72/9 = b

OpenStudy (anonymous):

yep b is simply a/h=72/9=8